OpenStudy (anonymous):

find two consecutive positive integers such that the sum of their squares is 61?????

OpenStudy (anonymous):

61+n+n+1 ???

OpenStudy (anonymous):

the 2 numbers can be represented by x and x+1 x^2+(x+1)^2=61

OpenStudy (anonymous):

30 & 31?

OpenStudy (anonymous):

the sum of their squares is 31, not the sum of the numbers

OpenStudy (anonymous):

x^3+2=61

OpenStudy (anonymous):

x^2+(x+1)^2=61 x^2+x^2+2x+1=61 2x^2+2x+1=61 2x^2+2x-60=0 x^2+x-30=0 (x+6)(x-5)=0 x=-6,5 you want the positive x here so the first number is 5, the second number is 6 5^2+6^2=61 25+36=61 61=61

OpenStudy (anonymous):

thank you